OpenStudy (anonymous):

The density of a solution of sucrose is 1.0678g/cc, and it is 17.00 percent by mass. What volume of solution in milliliters do you need to supply 41.0g of surcrose

OpenStudy (rushwr):

@Photon336

OpenStudy (photon336):

@Rushwr sorry cc what unit is that?

OpenStudy (rushwr):

cubic centimeter

OpenStudy (rushwr):

@Photon336 if the solution is 17% by mass then the other 83% should sucrose?

OpenStudy (rushwr):

be ^

OpenStudy (photon336):

\[\frac{ 1.0678g }{ cm ^{3} }\] @Rushwr I think that 17% by mass is it's 17% sucrose by mass @peachpi what do you think?

OpenStudy (anonymous):

that's what I think as well

OpenStudy (rushwr):

Then I guess the mass of solution is 83%

OpenStudy (photon336):

I think that can find % by volume if i'm correct from that

OpenStudy (anonymous):

so if there's 41 g of sucrose, then the mass of the whole solution is 0.17x=41 x=241.2 g ?

OpenStudy (photon336):

\[\frac{ 41 }{ x } = \frac{ 17 }{ 100 }\] x = 241.2 grams @peachpi I totally agree

OpenStudy (anonymous):

and then from that I think use the density to find volume

OpenStudy (rushwr):

yeah neh !

OpenStudy (anonymous):

thank you so much guys!!! @peachpi @Rushwr @Photon336

OpenStudy (rushwr):

I didn't do anythn though !!!!!! Sorry about that !

OpenStudy (photon336):

\[\frac{ 1.067g }{ cm ^{3} } = \frac{ 241.2g }{ x}\] I guess if we had this amount of grams not sure about this

OpenStudy (photon336):

no wait I think that's wrong that's the mass by volume that has to be converted to cm^3 but the conversion factor is 1:1

OpenStudy (anonymous):

I think you end up the same place though\[\rho=\frac{ m }{ V }\] \[V=\frac{ m }{ \rho }=\frac{241~g}{1.0678~g/cm^3}\]

OpenStudy (photon336):

interesting thanks @peachpi